carton contains 10 eggs, 3 of which are cracked. If you randomly select 5 of the eggs for hard boiling, what is the probability of the following events?
i. All of the cracked eggs are selected.
ii. None of the cracked eggs are selected.
iii. Two of the cracked eggs are selected.

Answer by stanbon(75887)

carton contains 10 eggs, 3 of which are cracked. If you randomly select 5 of the eggs for hard boiling, what is the probability of the following events?
---
Binomial Problem with n = 5 and p(cracked) = 0.3 ; p(not cracked) = 0.7
--------
i. All of the cracked eggs are selected.
P(x = 3) = 5C3(0.3)^3*0.7^2 = 0.1323
-------------------------------------------
ii. None of the cracked eggs are selected.
P(x = 0) = 5C0*0.3^0*0.7^5 = 0.1681
-------------------------------------------
iii. Two of the cracked eggs are selected.
P(x = 2) = 5C2(0.3)^2*0.7^3 = 0.3087
=========================
Cheers,
Stan H.
=========================